5) Fill in the following table: P(atm) 0.50 V(L) n(moles) 2 1 0.50 T(K) 150 298.15 0.75 22.4 5 10 300 1500 0.15 0.005
Answer #1

1)

Given:

P = 0.5 atm

n = 2 mol

T = 150.0 K

use:

P * V = n*R*T

0.5 atm * V = 2 mol* 0.08206 atm.L/mol.K * 150 K

V = 49.236 L

Answer: 49 L

2)

Given:

V = 22.4 L

n = 1 mol

T = 298.15 K

use:

P * V = n*R*T

P * 22.4 L = 1 mol* 0.08206 atm.L/mol.K * 298.15 K

P = 1.0922 atm

Answer: 1.09 atm

3)

Given:

P = 0.75 atm

V = 5.0 L

n = 0.50 mol

use:

P * V = n*R*T

0.75 atm * 5 L = 0.5 mol* 0.08206 atm.L/mol.K * T

T = 91.3965 K

Answer: 91.4 K

4)

Given:

P = 2.0 atm

V = 10.0 L

T = 300.0 K

find number of moles using:

P * V = n*R*T

2 atm * 10 L = n * 0.08206 atm.L/mol.K * 300 K

n = 0.8124 mol

Answer: 0.812 mol

5)

Given:

P = 0.15 atm

n = 0.005 mol

T = 1500.0 K

use:

P * V = n*R*T

0.15 atm * V = 0.005 mol* 0.08206 atm.L/mol.K * 1500 K

V = 4.103 L

Answer: 4.1 L